# What Would an LLM Do? Evaluating Policymaking Capabilities of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are increasingly being adopted in high-stakes domains. Their capacity to process vast amounts of unstructured data, explore flexible scenarios, and handle a diversity of contextual factors can make them uniquely suited to provide new insights for the complexity of social policymaking. This article evaluates whether LLMs’ are aligned with domain experts (and among themselves) to inform social policymaking on the subject of homelessness alleviation – a challenge affecting over 150 million people worldwide. We develop a novel benchmark comprised of decision scenarios with policy choices across four geographies (South Bend, USA; Barcelona, Spain; Johannesburg, South Africa; Macau SAR, China). The policies in scope are grounded in the conceptual framework of the Capability Approach for human development. We also present an automated pipeline that connects the benchmarked policies to an agent-based model, and we explore the social impact of the recommended policies through simulated social scenarios. The paper results reveal promising potential to leverage LLMs for social policy making. If responsible guardrails and contextual calibrations are introduced in collaboration with local domain experts, LLMs can provide humans with valuable insights, in the form of alternative policies at scale.
Introduction
@@
Homelessness, defined by the United Nations (UN) as “the lack of a stable, safe, and adequate housing” (Special Rapporteur on the right to adequate housing 2022), is a growing crisis affecting cities and nations worldwide. According to the UN, an estimated 1.6 billion people globally lack adequate housing, with 150 million completely homeless. The OECD reports that in most developed countries, homelessness has sharply increased since 2022 (OECD 2024), with the U.S. alone recording more than 653,000 unhoused individuals in a single night in January 2023 (Daniel Soucy and Hall 2024). Nearly 15.6 % of the population of Johannesburg lived on the streets in 2022 (Statistics South Africa 2025), while Barcelona’s homeless population has grown by 61% between 2008 and 2022 (Xarxa d’Atenció a Persones Sense Llar (2023) XAPSLL). Homelessness represents a severe deprivation of basic human security and dignity, undermining sustainable development goals, and tackling its challenges is a core requirement for creating equitable, resilient societies where no one is left behind (United Nations 2015).
@@
In response to these urgent social challenges, policymaking to alleviate homelessness should extend beyond material redistribution to confront a more profound need: ensuring unhoused individuals are treated as equals and fully integrated into the social fabric. This necessitates dismantling structural stigmatization and fostering genuine belonging—a task complicated by deeply entrenched societal biases, bureaucratic systems that often reduce people to “cases”, and the trauma-induced isolation experienced by those without stable shelter (Curto et al. 2025; Ranjit et al. 2024). As Wasserman and Clair (2009) underscore, homeless individuals frequently face “institutional humiliation” in shelters or services, exacerbating their marginalization. The World Bank’s Voices of the Poor further reveals that exclusion persists even after housing access, as societal rejection impedes employment, community participation, and psychological recovery (Narayan et al. 2000). Thus, effective homelessness mitigation policies must not only allocate resources but actively combat dehumanization, centering the agency and dignity of those affected, and recognize inclusion as both an ethical imperative and a pragmatic necessity.
@@
Given the multifaceted nature of homelessness and the difficulty of anticipating policy outcomes in dynamic social systems, computational approaches offer a promising avenue for navigating this complexity and enable prospective policy testing. Recent advances in computational social science have demonstrated the potential of agent-based models (ABMs) to inform homelessness policymaking through dynamic modeling of complex socio-economic systems. Aguilera et al. (2025) developed an innovative framework for human development with ABMs, representing homelessness as a multidimensional deprivation, where agents interact within policy-sensitive environments. However, significant challenges remain in modeling human behavior within social simulations, such as the inability of capturing latent psychological factors like values, needs and emotional states (Dignum 2021), and the difficulty of developing generalizable models due to the dramatic differences in local policies, social networks, and resource availability across geographies (Aguilera et al. 2025). To address this limitation, there is active research on connecting ABMs with LLMs, demonstrating promising results in exploratory studies (Anthis 2025).
@@
Simultaneously, we note that the increasing popularity of AI-driven tools like LLMs and their pervasive integration into societal systems make them increasingly likely to inform policy makers in various contexts, such as in finance (Febrian and Figueredo 2024) and law (Yun and Lee 2025). However, deploying LLMs in socially sensitive domains, such as homelessness policymaking, presents serious challenges. These include the amplification of societal biases (Agnew et al. 2024), generation of plausible but misleading outputs (“hallucinations”), a lack of genuine understanding of human suffering and social dynamics (Shanahan 2024), and inherent opacity that makes their reasoning difficult to audit. It is therefore imperative to proactively assess their risks and limitations in such high-stakes contexts.
@@
To address these concerns, we introduce a benchmark with decision scenarios involving policy choices aimed at reducing homelessness, to assess the alignment of LLMs with domain experts in four different parts of the world (South Bend Indiana, USA; Barcelona, Spain; Johannesburg, South Africa; Macau SAR, China). The selection of policies included in the benchmark is grounded in the conceptual framework of the Capability Approach for human development (CA) (Sen 1999; Nussbaum 2011; Robeyns 2017). Unlike models that systematically prioritize the material needs of individuals (Maslow 1943), the CA considers opportunities that human beings have to lead a meaningful life with dignity (Sen 1999). From this perspective, homelessness can be considered as a deprivation of central capabilities (Nussbaum 2011). We perform an extensive empirical investigation around LLMs’ choices and judgments on the benchmark, including comparing them with those offered by domain experts in the geographical locations.
@@
Finally, we propose a novel pipeline to automatically link policies with an ABM framework, intended to gauge the social impacts of LLM generated policies in a simulated social context. Specifically, this approach explores how policies inform agent behavior through an existing State-Action-Transition process (SAT) (Aguilera et al. 2024). Thus, the pipeline establishes a connection between algorithmic policy suggestion and simulated societal impact. Our work opens up new directions to leverage LLMs for supporting social policymaking in a manner that is scalable and non-invasive, where LLMs could be used to suggest new packages of policies aligned with the conceptual framework of the Capability Approach, and their impact could be examined via ABM simulations.
Related Work
The Capability Approach in Computational Frameworks
@@
Although Sen and Nussbaum’s work on the Capability Approach (Sen 1999; Nussbaum 2011) constitutes a cornerstone in human development studies and the conceptual framework that underpins the Sustainable Development Goals, its operationalization in the practice of social policy making still has much potential to unveil (Robeyns 2017). From a computational social practice perspective, only recently has the CA been integrated into computational methods. There is an incipient corpus that aims to integrate the conceptual framework of the CA in agent-based modeling (Aguilera et al. 2025; Chávez-Juárez and Krishnakumar 2021), but to the best of our knowledge, our work constitutes the first application of CA as an audit framework for generative AI in social policy.
Evaluating LLMs for Social Policymaking
@@
Recent research has explored the potential and limitations of LLMs in policymaking contexts. Jiao et al. (2025) systematically documented risks when deploying LLMs in high-stakes policy domains, highlighting issues of bias amplification and context blindness, which this paper aims to address. Work by Wei, Kumar, and Zhang (2025) provided evidence of how LLMs can perpetuate societal biases in resource allocation scenarios, raising critical questions about their suitability for welfare policymaking. The recurrent hallucinations in general problem solving (Xu, Jain, and Kankanhalli 2025) takes on particular significance in homelessness policymaking, where factual inaccuracies could directly impact the most vulnerable individuals.
@@
While frameworks exist for evaluating LLM outputs in healthcare decision making (Kanithi et al. 2024) and in legal domains (Li et al. 2024), homelessness policy presents unique challenges due to its intersection with complex socioeconomic factors, stigma dynamics, and diverse cultural contexts. Our work presents a novel framework to specifically assess the ability of LLMs to inform policy making on the topic of homelessness alleviation.
Integrating LLMs with Agent-Based Social Simulations
@@
The integration of LLMs with ABMs constitutes an emerging topic in computational social science (Chopra et al. 2024). Traditional ABMs have proved very helpful to inform social policy making in a non-invasive manner (Dignum 2021). LLMs have shown the potential to enrich the behavior of agents with reasoning-like abilities, memory, personality traits, and context sensitivity (Ferraro et al. 2025; Park et al. 2023; Horto 2023). Although conceptual frameworks have been presented to reliably connect LLMs with ABMs in socially sensitive frameworks (Ricci et al. 2024), the technical approaches still remain a challenge. In this work, we explore the potential impact of policies from our proposed benchmark in a simulated social context by adapting the behavior of agents accordingly in an agent-based modeling simulation context.
Methodology
@@
Figure 1: Methodology overview: we construct a benchmark via a structured prompting strategy grounded in the Capability Approach, prompt various LLMs to act as policymakers, and analyze their choices through comparison with human expert judgments and a study of their projected societal impact using a modular agent-based modeling pipeline.
Benchmark Design
We present the components of a novel benchmark that aims to evaluate whether LLMs are aligned with domain experts, within the locations in scope, to inform homelessness alleviation policymaking [^1] The benchmark will be made publicly available later..
Dataset Composition
@@
The benchmark covers four geographic contexts: Barcelona (Spain), Johannesburg (South Africa), South Bend (USA), Macau SAR (China), as well as a Universal context not tied to any particular city. The universal context is included to study the impact of context on choices. For each of the four specific locations, we generate 40 decision-making scenarios, while the universal category contains 10 scenarios. Each scenario frames a dilemma faced by a non–profit organisation (NPO) working to alleviate homelessness, e.g., how to allocate emergency housing funds, whether to prioritise medical outreach or job training, or how to intervene in the face of gentrification pressures. In total, the benchmark comprises 170 scenarios, each with 4 policy choices.
@@
Every scenario is presented as a structured JSON object with four fields: Scenario (a title summarising the dilemma), Context (one or more paragraphs providing background), Policy_Options (an array of four policy proposals) and Main_capability_restoration (a list of capabilities targeted by each policy). The context is written in narrative form and spans at least 80 words, capturing locally relevant demographics, political debates, economic constraints and historical factors. Each policy option is described in a paragraph of at least 35 words and is annotated with one or more of Martha Nussbaum’s ten central human capabilities (see Appendix A) [^2] All Appendices are in the supplementary material.. These capabilities, summarized in the literature as universal freedoms that societies ought to protect, range from the ability to live a full life and access healthcare to the freedom to develop practical reason and to participate in political decision making (Nussbaum 2012). Using this normative framework ensures that each policy speaks to a specific aspect of human flourishing rather than merely enumerating material resources.
Scenario Generation
@@
To seed the generation process, we first manually crafted a baseline scenario for Barcelona in consultation with domain experts. This scenario, rooted in current debates about housing scarcity (Xarxa d’Atenció a Persones Sense Llar (2023) XAPSLL), presented a homelessness dilemma alongside four plausible policy responses aiming to restore affected central capabilities. The manual scenario served both as a template for automated generation and as a qualitative anchor for calibrating LLM outputs.
@@
Subsequent scenarios were generated using a frontier LLM (GPT–4.1) instructed to act as a public policy expert with knowledge on urban development and the Capability Approach. We provided the LLM with a detailed project description, a list of the ten central capabilities, guidance on scenario length and tone, and examples of desirable output. The model was then prompted to produce exactly 40 non–redundant scenarios for each city (and 10 for the universal category). Each prompt emphasized that scenarios must (1) present a localized public policy dilemma related to homelessness, (2) be either grounded in a documented real situation or be a plausible fictional event, (3) include context paragraphs of sufficient length, (4) propose four diverse and non–redundant policy options, and (5) annotate each policy with its primary capability restoration. To avoid hallucinated or extreme situations, we explicitly instructed the model not to invent unlikely events (e.g., large scale violence in low–crime areas) and to “think like a local expert who has studied homelessness for many years”. Generation was performed separately for each location to ensure contextual fidelity.
@@
Generated scenarios were post–processed to remove redundancies and to correct factual inaccuracies. When the LLM produced incomplete or duplicate content, we regenerated the scenario with a revised prompt. We also verified that the distribution of capabilities across policy options was balanced so that no capability was systematically over-represented, and that the language was neutral and devoid of harmful stereotypes.
Policy Annotation and Ethical Grounding
Beyond the narrative context, every policy option in the benchmark is linked to one or more human capabilities. These annotations (explicitly linking to capabilities) serve two roles: (i) they can guide evaluators (human or model) toward understanding the ethical aspirations behind each intervention, and (ii) they provide a means for measuring whether LLM recommendations favour certain types of capabilities over others. Annotating the policies in the framework of the Capability Approch involved identifying the core goal of each policy—whether it primarily protects life and physical health, fosters social belonging, supports practical agency, or promotes other types of individual freedom—and mapping that to the relevant capability.
Human Expert Annotation
To create a gold-standard baseline against which LLMs can be compared, we engaged domain experts from three out of the four cities in the benchmark – Barcelona, Johannesburg, and South Bend. Each expert is a practitioner or scholar with contextualized knowhow on homelessness alleviation landscape in their city and works in close collaboration with local non–profits. We provided them with 10 out of the 40 scenarios relevant to their location as well as the 10 universal scenarios, asking them to prioritise the four policies in each scenario from 1 (most preferred) to 4 (least preferred). Experts were instructed that there was no single correct answer and that they should rank options based on their topic and local knowledge, as well as their ethical considerations. Through this process, we obtained three sets of human-derived rankings—one for each city—that serve as the benchmark’s reference recommendations.
Model Evaluation Tasks

### v2

# What Would an LLM Do? Evaluating Large Language Models for Policymaking to Alleviate Homelessness

@@ Abstract
Large language models (LLMs) are increasingly being adopted in high-stakes domains. Their potential to encode evolving social contexts and to generate plausible scenarios position them as promising tools in social policymaking. This article evaluates whether LLMs are aligned with domain experts (and among themselves) on policy recommendations to alleviate homelessness – a challenge affecting over 150 million people worldwide. We develop a novel benchmark comprised of decision scenarios across four cities, with policy choices that are grounded in the conceptual framework of the Capability Approach for human development. We also present an automated pipeline that connects the policies to an agent-based model in one location, and compare the social impact of the policies recommended by LLMs to those recommended by experts. Our exploratory analysis reveals variation across LLMs in their policy recommendations compared to local experts, yet suggests potential benefits of the use of LLMs to provide insights for policymaking, if paired with responsible guardrails, contextual calibrations, and local domain expertise. Our work operationalizes the Capability Approach in a computational framework and provides new insights on homelessness alleviation policymaking with a focus on human dignity.
1 Introduction
@@
Homelessness, defined by the United Nations (UN) as “the lack of a stable, safe, and adequate housing” Special Rapporteur on the right to adequate housing (2022), is a growing crisis affecting cities and nations worldwide. According to the UN, an estimated 1.6 billion people globally lack adequate housing, with 150 million completely homeless. The OECD reports that in most developed countries, homelessness has sharply increased since 2022 OECD (2024), with the U.S. alone recording more than 653,000 unhoused individuals in a single night in January 2023 Daniel Soucy and Hall (2024). Homelessness represents a severe deprivation of basic human security and dignity, undermining sustainable development goals, and tackling these challenges constitutes a core requirement for creating equitable, resilient societies where no one is left behind United Nations (2015).
@@
In response to these urgent social challenges, policymaking to alleviate homelessness needs to extend beyond material redistribution to confront a more profound requirement: ensuring unhoused individuals are treated as equals and fully integrated into the social fabric. This necessitates dismantling structural stigmatization and fostering genuine belonging – a task complicated by deeply entrenched societal biases, bureaucratic systems that often reduce people to “cases”, and the trauma-induced isolation experienced by those without stable shelter Curto et al. (2025); Ranjit et al. (2024).
@@
Effective homelessness mitigation policies must not only allocate material resources but actively combat dehumanization and societal rejection Takahashi (1998); Narayan et al. (2000); Wasserman and Clair (2009), centering the agency and dignity of those affected, and recognize inclusion as both an ethical imperative and a pragmatic necessity. Given the multifaceted nature of homelessness and the difficulty in anticipating policy outcomes in dynamic social systems, computational approaches offer a promising avenue to navigate this complexity and enable prospective policy testing. Recent advances in computational social science have demonstrated the potential of agent-based models (ABMs) to inform homelessness policymaking, representing homelessness as a multidimensional deprivation Aguilera et al. (2025). However, significant challenges remain in modeling human behavior within social simulations, such as the complexity to capture latent psychological factors and social values  Dignum (2021), as well as the difficulty around developing generalizable models that align with different local policies, evolving social contexts, and resource availability across geographies Aguilera et al. (2025)
@@
In this context, the increasingly widespread integration of LLMs in societal systems, their capacity to process vast amounts of unstructured data, explore flexible role-play scenarios, and handle a diversity of evolving contextual factors can make them uniquely suited to provide new insights for decision making in complex scenarios Wang (2024); Gao et al. (2023); Guo et al. (2024). Despite these capabilities, deploying LLMs in socially sensitive domains, such as homelessness policymaking, presents severe challenges. These include the amplification of societal biases Agnew et al. (2024), generation of plausible but misleading outputs (“hallucinations”), a lack of genuine understanding of human suffering and social dynamics Shanahan (2024), and inherent opacity that makes their reasoning difficult to audit. It is therefore imperative to proactively assess their risks and limitations in such high-stakes contexts.
@@
This paper aims to address these limitations by comparing recommendations made by LLMs with those provided by domain experts. We introduce a benchmark with decision scenarios that contain policy choices to reduce homelessness in four different parts of the world (South Bend Indiana, USA; Barcelona, Spain; Johannesburg, South Africa; Macau SAR, China). The selection of policies included in the benchmark is grounded in the conceptual framework of the Capability Approach for human development (CA) Sen (1999); Nussbaum (2011); Robeyns (2017). Unlike models that systematically prioritize the material needs of individuals Maslow (1943), the CA considers the opportunities that human beings have to lead a meaningful life with dignity Sen (1999). From this perspective, homelessness can be considered as a deprivation of central capabilities Nussbaum (2011). In our benchmark, the range of policies offered in the different scenarios illustrate the restoration of CA’s central capabilities for population in a situation of homelessness. We perform an empirical evaluation around LLMs’ choices and judgments on the benchmark, comparing them with those offered by the domain experts in each geographical location.
@@
Then, we propose a novel pipeline to automatically link policies with an ABM framework, intended to gauge the social impacts of LLM generated policies, versus those proposed by domain experts, in a simulated social context. Specifically, this approach explores how policies inform agent behavior through an existing State-Action-Transition process (SAT) Aguilera et al. (2024). Our work opens up new directions to leverage LLMs for supporting social policymaking in a manner that is scalable and non-invasive, where LLMs could be used to suggest alternative policies aligned with the conceptual framework of the CA, and their impact could be examined via ABM simulations. The paper operationalizes the Capability Approach for human development in a computational framework, suggesting policies that go beyond covering the basic material needs of the persons with no home, with a focus on the whole social fabric and human dignity. The work is requested and conducted in close collaboration with domain experts and non-profit organizations specialized in alleviating homelessness in the four cities in scope. The domain experts participated in the definition of the scope and research approach. Multi-disciplinary domain experts defined the baseline scenario and conducted policy rankings in the benchmark scenarios. Insights from our work are being leveraged to inform ongoing policymaking discussions in the four locations in scope.
2 Related Work
2.1 The Capability Approach
@@
Sen and Nussbaum’s work on the Capability Approach constitutes a cornerstone in human development studies and as a conceptual framework underpinning the Sustainable Development Goals. And its operationalization in the practice of social policy making still has much potential to unveil Robeyns (2017). From a computational social practice perspective, only recently has the CA been integrated into methods such as agent-based modeling Aguilera et al. (2025); Chávez-Juárez and Krishnakumar (2021) To the best of our knowledge, our work constitutes the first application of CA as a framework for generative AI in social policy.
2.2 Evaluating LLMs for Social Policymaking
@@
Recent research has explored the potential and limitations of LLMs in policymaking contexts. Jiao et al. (2025) systematically documented risks when deploying LLMs in high-stakes policy domains, highlighting issues of bias amplification and context blindness, which this paper aims to address. Work by Wei et al. (2025) provided evidence of how LLMs can perpetuate societal biases in resource allocation scenarios, raising critical questions about their suitability for welfare policymaking. Recurrent hallucinations in general problem solving Xu et al. (2025) takes on particular significance in homelessness policymaking, where factual inaccuracies could directly impact the most vulnerable individuals.
@@
While frameworks exist for evaluating LLM outputs in healthcare decision making Kanithi et al. (2024) and in legal domains Li et al. (2024), homelessness alleviation policy presents unique challenges due to its intersection with complex socioeconomic factors and social stigmatization, which have different nuisances across geographies. Our work presents a novel framework to specifically assess the ability of LLMs to inform policy making on the topic of homelessness alleviation in four distinct geographic locations.
2.3 Integrating LLMs with Agent-Based Social Simulations
@@
The integration of LLMs with ABMs in social simulations constitutes an emerging topic in computational social science Chopra et al. (2024); Guo et al. (2024); Gao et al. (2024); Jin et al. (2025). Traditional ABMs have proven to be very helpful to inform social policy making in a non-invasive manner Dignum (2021). LLMs have shown the potential to enrich and diversify the behavior of agents with reasoning-like abilities, memory, personality traits, and context sensitivity Ferraro et al. (2025); Park et al. (2023); Horto (2023). Although conceptual frameworks have been presented to connect LLMs with ABMs in socially sensitive frameworks Ricci et al. (2024), the technical complexities remain a challenge and there is still a question about the reliability of recommendations provided by LLMs as inputs for policymaking. In this work, we present a pipeline that connects the policies recommended by LLMs, and we evaluate their impact on homelessness alleviation in an ABM framework.
3 Methodology
@@
Figure 1: Methodology overview: We construct a benchmark via a structured prompting strategy grounded in the Capability Approach for human development (CA) prompt various LLMs to act as policymakers, analyze their choices through comparison with human expert recommendations, and conduct an evaluation of their projected societal impact using a modular agent-based modeling pipeline.
3.1 Benchmark Design
We present the components of a novel benchmark that aims to evaluate whether LLMs are aligned with domain experts working with specialized organizations, within the locations in scope, to inform homelessness alleviation policymaking [^1] We will make the benchmark publicly available..
3.1.1 Dataset Composition
@@
The benchmark covers four geographic contexts: Barcelona (Spain), Johannesburg (South Africa), South Bend (USA), Macau SAR (China), as well as a Universal context not tied to any particular city. The universal context is included to study the impact of context on choices. For each of the four specific locations, we generate 40 decision-making scenarios, while the universal category contains 10 scenarios. Each scenario frames a dilemma faced by a non–profit organization working to alleviate homelessness, e.g., how to allocate emergency housing funds, whether to prioritize medical outreach or job training, etc. In total, the benchmark comprises 170 scenarios, each with 4 choices.
@@
Every scenario is presented as a structured JSON object with four fields: Scenario (a title summarising the dilemma), Context (one or more paragraphs providing background), Policy_Options (an array of four policy proposals) and Main_capability_restoration (a list of central capabilities targeted by each policy). The context is written in narrative form and spans at least 80 words, capturing locally relevant demographics, political debates, economic constraints and historical factors. Each policy option is described in a paragraph of at least 35 words and is annotated with one or more of Martha Nussbaum’s ten central human capabilities (see Appendix A) [^2] Appendices are in the supplementary material., which are universal freedoms that societies ought to protect – ranging from the ability to live a full life and access healthcare to the freedom to develop practical reason and to participate in political decision making Nussbaum (2012). Using this normative framework ensures that each policy speaks to a specific aspect of human flourishing rather than merely enumerating material resources.
3.1.2 Scenario Generation
@@
To seed the generation process, we first manually crafted a baseline scenario for Barcelona in consultation with domain experts. This scenario, rooted in current debates about housing in the city Xarxa d’Atenció a Persones Sense Llar (XAPSLL) (2023), presented a homelessness dilemma alongside four plausible policy responses aiming to restore affected central capabilities. The manual scenario served both as a template for automated generation and as a qualitative anchor for calibrating LLM outputs.
@@
Subsequent scenarios were generated using a frontier LLM (GPT–4.1) instructed to act as a public policy expert with knowledge on urban development and the Capability Approach. We provided the LLM with a detailed project description, a list of the ten central capabilities, guidance on scenario length and tone, and examples of desirable output. The model was prompted to produce exactly 40 non–redundant scenarios for each city (and 10 for the universal category). Each prompt emphasized that scenarios must (1) present a localized public policy dilemma related to homelessness, (2) be either grounded in a documented real situation or be a plausible fictional event, (3) include context paragraphs of sufficient length, (4) propose four diverse and non–redundant policy options, and (5) annotate each policy with its primary capability restoration. To avoid extreme situations, we explicitly instructed the model not to invent unlikely events (e.g., large scale violence in low–crime areas) and to “think like a local expert who has studied homelessness for many years”. Generation was performed separately for each location and scenarios were manually reviewed by our team of researchers and multi-disciplinary domain experts to ensure realism and eliminate incoherent answers.
@@
Generated scenarios were post-processed to remove redundancies and inaccuracies. When the LLM produced incomplete or duplicate content, we regenerated the scenario with a revised prompt. We also verified that the distribution of capabilities across policies was balanced, and that the language was neutral and devoid of harmful stereotypes.
3.1.3 Policy Annotation and Ethical Grounding
Beyond the narrative context, every policy option in the benchmark is linked to one or more human capabilities. These tags serve two roles: (i) they can guide evaluators (human or model) toward understanding the ethical aspirations behind each intervention, and (ii) they provide a means for measuring whether LLM recommendations favor certain types of capabilities over others. Annotating the policies in the framework of the Capability Approach involved identifying the core goal of each policy –whether it primarily protects life and physical health, fosters social belonging, supports practical agency, or promotes other types of individual freedom – and mapping that to the relevant capability.
3.1.4 Expert Annotation
To create a human baseline against which LLMs can be compared, we engaged two domain experts (one primary expert and a secondary one) from three out of the four cities in the benchmark – Barcelona, Johannesburg, and South Bend. Experts are multi-disciplinary scholars working on homelessness alleviation in non-profit organizations or in close collaboration with local non-profits in the cities in scope.
We provided experts with 10 out of the 40 scenarios relevant to their location as well as the 10 universal scenarios, asking them to prioritize the four policies in each scenario from 1 (most preferred) to 4 (least preferred). Experts were instructed that there was no single correct answer and that they should rank options based on their topic and local knowledge, as well as their ethical considerations. Through this process, we obtained six sets of human-derived rankings – two for each city – that serve as the benchmark’s human reference recommendations.
3.1.5 Model Evaluation Tasks